OpenStudy (anonymous):

Which of the following values of a will result in a linear equation with no solutions? 8x+a=8x−7

OpenStudy (texaschic101):

are there answer choices ?

OpenStudy (texaschic101):

a could be -7 and it would make it where there is no solution 8x + 7 = 8x - 7 8x - 8x = -7 - 7 0 = -14 (incorrect...no solution)